# Solutions to multiplicative Diophantine equations in orthotopes

### v1

@@ 1.1. Set-up
Our paper consists of five main sections. The main results are stated in §2, with Theorem 2.1 being the most general result of this work, and Theorem 2.2 being, under some additional conditions, the strongest result. Then, Theorems 2.3 and 2.4 demonstrate how our main results can be used to derive explicit asymptotic formulæ for the power moments of restricted divisor functions. These theorems are then applied to other number theoretical problems, as discussed in §2.3.
@@ 1.3. Previous results
Before the work of de la Bretèche, Kurlberg, and Shparlinski [9], the only known results for this problem were for the smaller cases, sometimes with additional coprimality conditions. For example, for $m=\ell=2$, Tolev [41] gave the first known results on the asymptotics of $\mathsf{E}(\mathbf{1}_{2}\oplus(2);H^{\mathbf{1}_{(2,1)}})$. For $m=\ell=3$, de la Bretèche [6], Fouvry [19], and Heath-Brown and Moroz [26] gave related results on the number of integer solutions of this equation.
@@ 1.4. Number theoretical applications
Indeed, it was shown by the first-named author [1], with respect to our notation, that
(1.12) $\sum_{n}\tau(n;H)^{2}=\mathsf{E}(\mathbf{1}_{(2,2)};H^{\mathbf{1}_{(2,2)}})=\dfrac{12}{\pi^{2}}H^{2}\log H+O(H^{2}),$
which was previously shown by Mastrostefano[29] with a non-explicit leading term. Moreover, a weighted version of (1.12) with a non-explicit secondary main term was proven by Heath-Brown [25]. In particular, for a certain class of smooth weights $w:{\mathbb{Z}}^{4}\to{\mathbb{R}}_{+}$, he showed that
@@
for some constants $C_{j}(w)$, depending at most on $w$, where $C_{1}(w)>0$. Now, in Corollary 2.5 below, we derive a non-explicit secondary main term and get a power saving for the error term in (1.12). Note that $\tau(n;H)=0$ whenever $n>H^{2}$, and thus the left-hand side of (1.12) is well-defined.
Now, the function $\tau(n;H)$ first appeared in Truelsen [42], under a slightly different notation, who used this function in relation to the problem of determining the irrational numbers $\alpha$ for which the pair correlation for the fractional parts of $n^{2}\alpha$ is Poissonian. The same function also appears in Mastrostefano [28], who used it to classify maximal size product sets of random subsets of ${\mathbb{Z}}$. In this article, we discuss a generalisation of this function, the generalised restricted divisor function
@@ 2.2. Applications to restricted divisor functions
$\sum_{n}\tau_{m}(n;H^{\mathbf{c}})^{k}=q_{m,k,\mathbf{c}}H^{\|\mathbf{c}\|}(\log H)^{m^{k}-(m-1)k-1}+H^{\|\mathbf{c}\|}P_{m,k,\mathbf{c}}(\log H)+o(H^{m-\vartheta_{m,k,\mathbf{c}}}),$
@@
2.3. Applications to counting singular matrices
A special application of Theorem 2.3 is the following corollary, which improves (1.12), and consequently a handful of other results in the recent work [1] of the first-named author.
Corollary 2.5.
There exists absolute constants $A$ and $\vartheta>0$ such that
$\sum_{n}\tau(n;H)^{2}=\frac{12}{\pi^{2}}H^{2}\log H+AH^{2}+o(H^{2-\vartheta})\quad\text{as}\quad H\to\infty.$
Proof.
The sum of interest pertains to Theorem 2.3 in the case where $m=k=2$ and $\mathbf{c}=\mathbf{1}_{2}$. Thus, by virtue of Theorem 2.3, it suffices to show that $q_{2,2,\mathbf{1}_{2}}=12/\pi^{2}$. Firstly, note that $V_{2,2,\mathbf{1}_{2}}$ is the $12^{2}-2(2-1)-1=$ 1 dimensional volume of a region embedded in a $42^{2}=$ 4 dimensional space, that is, the length of a line segment in ${\mathbb{R}}^{4}$. Therefore
and consequently
$q_{2,2,\mathbf{1}_{2}}=2\prod_{p}(1-p^{-1})^{4}\sum_{n\geqslant 0}\frac{(n+1)^{2}}{p^{n}}=2\prod_{p}(1-p^{-2})=\frac{2}{\zeta(2)},$
[…]
We now describe applications of Corollary 2.5 to some other number theoretical problems. First, consider the problem of counting the number of elements in the set ${\mathcal{D}}_{2}(H;\Delta)$ of $2\times 2$ integer matrices of determinate $\Delta$ whose entries are integers of absolute value not exceeding $H$.
When $\Delta\neq 0$, asymptotic formulæ for #$\#{\mathcal{D}}_{2}(H;\Delta)$ have been given in [1, 20, 23] for a large range of $\Delta$ relative to $H$. For $\Delta=0$, the first-named author [1, Section 1.1] proved the estimate
#$\#{\mathcal{D}}_{2}(H;0)=\dfrac{96}{\pi^{2}}H^{2}\log H+O(H^{2}).$
We now state the improvement of this result by using Corollary 2.5.
Corollary 2.6.
As $H\to\infty$, we have
#$\#{\mathcal{D}}_{2}(H;0)=\dfrac{96}{\pi^{2}}H^{2}\log H+(8A+16)H^{2}+O(H^{2-\min{(\vartheta,1)}}),$
with $A$ and $\vartheta$ as in Corollary 2.5.
Proof.
(2.7) $ad=bc,$
with $-H\leq a,b,c,d\leq H$.
$(4H+1)^{2}=16H^{2}+8H+1$
solutions to (2.7).
We may now assume $abcd\neq 0$. By considering all possible signs of $a$, $b$, $c$ and $d$, similar to [1, Section 3], we see that
#$\#{\mathcal{D}}_{2}(H;0)=8\sum_{n}\tau(n;H)^{2}+16H^{2}+O(H),$
[…]
Another corollary of our result concerns the problem of counting the number of $2\times 2$ singular matrices whose entries are Egyptian fractions (or unit fractions) of a bounded height, first introduced in [2]. Following the notation in [2], we consider the problem of counting the cardinality of
The first-named author, along with Kuperberg, Ostafe and Shparlinski [2, Theorem 2.5], proved that
By employing a nearly identical argument as in the proof of Corollary 2.6, we have the following asymptotical formula for #$\#{\mathcal{D}}_{2}({\mathbb{Z}}^{-1};H,0)$.
Corollary 2.7.
As $H\to\infty$, we have
#$\#{\mathcal{D}}_{2}({\mathbb{Z}}^{-1};H,0)=\dfrac{96}{\pi^{2}}H^{2}\log H+8AH^{2}+O(H^{2-\vartheta}),$
with $A$ and $\vartheta$ defined in Corollary 2.5.
@@ 5.1. Non-explicit results
This is an example of where the lower bound (1.6) is precisely of the correct order. Note that applying Theorem 2.2 directly to improve (5.3) is not possible in this case, due to unavailability of a vector $\mathbf{a}$ satisfying the condition (2.3) of Theorem 2.2. The condition (2.3) itself is necessary in order for the condition (3.4) of Lemma 3.2 be fulfilled. This occurrence happens for most choices of $\bm{\gamma}$ that are not of the form $\bm{\gamma}=\mathbf{1}_{\mathbf{m}}$. In addition, for many choices of $\bm{\gamma}$, we have $\min(\|\mathbf{b}_{1}\|,\ldots,\|\mathbf{b}_{k}\|)>\mu_{\bm{\gamma},\mathbf{b}}$, in which case, the trivial bounds (1.4) and (1.6) differ by a factor greater than $O_{\varepsilon}(H^{\varepsilon})$.
@@ 5.2. Imposing further restrictions
for some polynomial $T$ of degree not exceeding $\kappa_{\bm{\gamma}}(\mathbf{a})-1$. For example, taking ${\mathcal{N}}$ to be the set of four-tuples of square-free numbers and appealing to Corollary 2.5, we see that the number of singular $2\times 2$ matrices with positive square-free entries not exceeding $H$ is asymptotic to
@@
where ${\mathcal{M}}\subset{\mathbb{N}}$ denotes the set of square-free natural numbers. Similarly, we can use Corollary 2.5 and (5.4) to show that, for any natural number $q$, we have
@@
with an error term similar to that in Corollary 2.5. More generally, we can use (5.4) along with (4.4) to count the number of integer solutions to the multiplicative Diophantine equation (1.2) where only select ordinates are restricted. We simply offer the above two examples as a way of illustrating to the reader how this procedure can easily be followed by appealing to Theorem 2.2.
We note that our method does not, unfortunately, allow us to count solutions to (1.2) over the Piatetski-Shapiro or Beatty numbers, as in these cases ${\mathcal{N}}$ is not closed under the Hadamard product. This is also the main obstacle for considering the case where we restrict to solutions lying in fixed residue classes.
@@ Acknowledgements
The authors would like to thank Alina Ostafe, Igor E. Shparlinski, and Liangyi Zhao for supports and suggestions during the preparation of this work. The authors would also like to thank Régis de la Bretèche, as well as Marc Munsch and Igor E. Shparlinski, for helpful correspondences regarding the articles [7], and [32], respectively. M.A. is supported by Australian Research Council Grants DP230100530 and a UNSW Tuition Fee Scholarship. C.C.C. is supported by the Australian Government and UNSW through RTP and UPA Scholarships, respectively.

### v2

@@ 1.1. Set-up
Our paper consists of five main sections. The main results are stated in §2, with Theorem 2.1 being the most general result of this work, and Theorem 2.2 being, under some additional conditions, the strongest result. Then, Theorems 2.3 and 2.4 demonstrate how our main results can be used to derive explicit asymptotic formulæ for the power moments of restricted divisor functions.
@@ 1.3. Previous results
Before the work of de la Bretèche, Kurlberg, and Shparlinski [10], the only known results for this problem were for the smaller cases, sometimes with additional coprimality conditions. For example, for $m=2$, Tolev [46] gave the first known results on the asymptotics of $\mathsf{E}(\mathbf{1}_{2}\oplus(\ell);H^{\mathbf{1}_{(2,1)}})$. For $m=\ell=3$, de la Bretèche [7], Fouvry [20], and Heath-Brown and Moroz [28] gave related results on the number of integer solutions of this equation, with additional coprimality conditions.
@@ 1.4. Number theoretical applications
Indeed, it was shown by Ayyad, Cochrane, and Zheng [3], with respect to our notation, that
(1.12) $\sum_{n}\tau(n;H)^{2}=\mathsf{E}(\mathbf{1}_{(2,2)};H^{\mathbf{1}_{(2,2)}})=\frac{2}{\zeta(2)}H^{2}\log H+AH^{2}+O\big{(}H^{19/13}(\log H)^{7/13}\big{)},$
where
$A=\frac{12\gamma-6\zeta^{\prime}(2)/\zeta(2)-3-6\zeta(2)}{3\zeta(2)}.$
A slightly weaker version of the above has recently been proven by Mastrostefano [33] and the first-named author [1] using vastly different techniques. Note that $\tau(n;H)=0$ whenever $n>H^{2}$, and thus the left-hand side of (1.12) is well-defined.
Counting instead in terms of the Euclidean norm, Louboutin and Munsch [31] showed that
#$\#\{\mathbf{x}\in\mathfrak{S}(X):x_{1}x_{2}=x_{3}x_{4}\}=\tfrac{3}{4}X^{2}\log X+O(X^{2}),$
where $\mathfrak{S}(X)=\{\mathbf{x}\in{\mathbb{N}}^{4}:\|\mathbf{x}\|_{2}\leqslant X\}$. Moreover, for a certain class of smooth weights $w:{\mathbb{Z}}^{4}\to{\mathbb{R}}_{+}$, Heath-Brown [27] showed that
@@
for some constants $C_{j}(w)$, depending at most on $w$, where $C_{1}(w)>0$ can be explicitly calculated.
Now, the function $\tau(n;H)$ first appeared in Truelsen [47], under a slightly different notation, who used this function in relation to the problem of determining the irrational numbers $\alpha$ for which the pair correlation for the fractional parts of $n^{2}\alpha$ is Poissonian. Truelsen [47] and Shparlinski [43] computed the sum of this function over a short arithmetic progression. The same function also appears in Mastrostefano [32], who used it to classify maximal size product sets of random subsets of ${\mathbb{Z}}$. In this article, we discuss a generalisation of this function, the generalised restricted divisor function
@@ 2.2. Applications to restricted divisor functions
$\sum_{n}\tau_{m}(n;H^{\mathbf{c}})^{k}=q_{m,k,\mathbf{c}}H^{\|\mathbf{c}\|}(\log H)^{m^{k}-(m-1)k-1}+H^{\|\mathbf{c}\|}P_{m,k,\mathbf{c}}(\log H)+o(H^{\|\mathbf{c}\|-\vartheta_{m,k,\mathbf{c}}}),$
@@ 5.1. Non-explicit results
This is an example of where the lower bound (1.6) is precisely of the correct order. Note that applying Theorem 2.2 to improve (5.3) is not directly possible in this case, due to unavailability of a vector $\mathbf{a}$ satisfying the condition (2.3) of Theorem 2.2. The condition (2.3) itself is necessary in order for the condition (3.4) of Lemma 3.2 be fulfilled. This occurrence happens for most choices of $\bm{\gamma}$ that are not of the form $\bm{\gamma}=\mathbf{1}_{\mathbf{m}}$. In addition, for many choices of $\bm{\gamma}$, we have $\min(\|\mathbf{b}_{1}\|,\ldots,\|\mathbf{b}_{k}\|)>\mu_{\bm{\gamma},\mathbf{b}}$, in which case, the trivial bounds (1.4) and (1.6) differ by a factor greater than $O_{\varepsilon}(H^{\varepsilon})$.
@@ 5.2. Imposing further restrictions
for some polynomial $T$ of degree not exceeding $\kappa_{\bm{\gamma}}(\mathbf{a})-1$. For example, taking ${\mathcal{N}}$ to be the set of four-tuples of square-free numbers and appealing to (5.8), we see that the number of singular $2\times 2$ matrices with positive square-free entries not exceeding $H$ is asymptotic to
@@
where ${\mathcal{M}}\subset{\mathbb{N}}$ denotes the set of square-free natural numbers. Similarly, we can use (5.8) and (5.4) to show that, for any natural number $q$, we have
@@
with an error term similar to that in (5.8).
More generally, we can use (5.4) along with (4.4) to count the number of integer solutions to the multiplicative Diophantine equation (1.2) where only select ordinates are restricted. We simply offer the above two examples as a way of illustrating to the reader how this procedure can easily be followed by appealing to Theorem 2.2.
We note that our method does not, unfortunately, allow us to count solutions to (1.2) over the Piatetski-Shapiro or Beatty numbers, as in these cases ${\mathcal{N}}$ is not closed under the Hadamard product. This is also the main obstacle for considering the case where we restrict to solutions lying in fixed residue classes. A possible exception of this case is when the class of residues are closed under Hadamard product, such as when we consider the set of quadratic or cubic residues modulo $\mathbf{q}$ for some integer vector $\mathbf{q}$, or when all solutions are $1$ modulo $\mathbf{q}$.
@@ 5.3. Other future directions
Yet another adaptation of our method can be used to count the number of solutions to equations of the form
$\alpha_{1}\mathbf{x}_{1}^{\bm{\gamma}_{1}}=\cdots=\alpha_{k}\mathbf{x}_{k}^{\bm{\gamma}_{k}}\pmod{K},$
for a fixed $K$, in a parallelpiped centered at the origin. The simplest case, $x_{1}x_{2}\equiv x_{3}x_{4}\pmod{p}$ when $k=p$ is a prime, was studied explicitly by Ayyad, Cochrane and Zheng [3], Garaev and Garcia [24] and Kerr [30] in an arbitrary box $B$ with character sums. We note that their method gives the best current result for counting solutions to the equation $x_{1}x_{2}=x_{3}x_{4}$ in a box, as stated in (1.12).
@@
5.4. Applications to counting singular matrices
Consider the sum of interest in Theorem 2.3 in the case where $m=k=2$ and $\mathbf{c}=\mathbf{1}_{2}$. Note that $V_{2,2,\mathbf{1}_{2}}$ is the $12^{2}-2(2-1)-1=$ 1 dimensional volume of a region embedded in a $42^{2}=$ 4 dimensional space, that is, the length of a line segment in ${\mathbb{R}}^{4}$. Therefore
and consequently
$q_{2,2,\mathbf{1}_{2}}=2\prod_{p}(1-p^{-1})^{4}\sum_{n\geqslant 0}\frac{(n+1)^{2}}{p^{n}}=2\prod_{p}(1-p^{-2})=\frac{2}{\zeta(2)}.$
This allows us to show that there exists a constant $\vartheta>0$ such that
(5.8) $\sum_{n}\tau(n;H)^{2}=\frac{2}{\zeta(2)}H^{2}\log H+AH^{2}+o(H^{2-\vartheta})$
for some constant $A$. Of course, the constant $A$ is the same as the one given in (1.12), which supersedes (5.8) in that the power-saving is explicit. Indeed, for the smaller cases such as this, elementary methods tend to give stronger results than the more general machinery of de la Bretèche, which is best-suited to the larger systems of variables considered in this article.
We now note that (5.8), and thus (1.12), improve on a key lemma of the first-named author [1] used for counting the number of elements in the set ${\mathcal{D}}_{2}(H;\Delta)$ of $2\times 2$ integer matrices of determinate $\Delta$ whose entries are integers of absolute value not exceeding $H$. Asymptotic formulæ for #$\#{\mathcal{D}}_{2}(H;\Delta)$ when $\Delta\neq 0$ have been given in [1, 21, 25] for a large range of $\Delta$ relative to $H$. For $\Delta=0$, the first-named author [1] showed that
#$\#{\mathcal{D}}_{2}(H;0)=\dfrac{96}{\pi^{2}}H^{2}\log H+O(H^{2}).$
We may now use (1.12) to improve this to the following.
Proposition 5.2.
As $H\to\infty$, we have
#$\#{\mathcal{D}}_{2}(H;0)=\dfrac{96}{\pi^{2}}H^{2}\log H+(8A+16)H^{2}+O(H^{19/13}(\log H)^{7/13}),$
where $A$ is as in (1.12).
Proof.
(5.9) $ad=bc,$
with $-H\leq a,b,c,d\leq H$.
$(4H+1)^{2}=16H^{2}+8H+1$
solutions to (5.9).
We may now assume $abcd\neq 0$. By considering all possible signs of $a$, $b$, $c$ and $d$, similar to as in §3 of [1], we see that
#$\#{\mathcal{D}}_{2}(H;0)=8\sum_{n}\tau(n;H)^{2}+16H^{2}+O(H),$
[…]
Another consequence of (1.12) concerns the problem of counting the number of $2\times 2$ singular matrices whose entries are Egyptian fractions (or unit fractions) of a bounded height, first introduced in [2]. Following the notation in [2], we consider the problem of counting the cardinality of
The first-named author, along with Kuperberg, Ostafe and Shparlinski [2], proved that
By employing a nearly identical argument as in the proof of Proposition 5.2, we have the following asymptotical formula for #$\#{\mathcal{D}}_{2}({\mathbb{Z}}^{-1};H,0)$.
Proposition 5.3.
As $H\to\infty$, we have
#$\#{\mathcal{D}}_{2}({\mathbb{Z}}^{-1};H,0)=\dfrac{96}{\pi^{2}}H^{2}\log H+8AH^{2}+O(H^{19/13}(\log H)^{7/13}),$
where $A$ is as in (1.12).
@@ Acknowledgements
The authors would like to thank Alina Ostafe, Igor E. Shparlinski, and Liangyi Zhao for supports and suggestions during the preparation of this work. The authors would also like to thank Régis de la Bretèche, as well as Marc Munsch and Igor E. Shparlinski, for helpful correspondences regarding their articles [8], and [36], respectively. The authors would again like to thank Marc Munsch for bringing the articles [31, 3] to their attention. M.A. is supported by Australian Research Council Grants DP230100530 and a UNSW Tuition Fee Scholarship. C.C.C. is supported by the Australian Government and UNSW through RTP and UPA Scholarships, respectively.